Alice deposits 1788 € for 3 months in a bank account which has an annual interest rate of 2.4 % (p.a.). Calculate the net accrued interest when the tax at source rate is 30 %.



1
Expert's answer
2022-02-03T18:08:03-0500

Solution:

"P=1788 \ €, t=3\ months, r=2.4\ \%\ p.a
\\ I=\dfrac{Prt}{100}=\dfrac{1788\times3\times2.4}{100\times 12}=10.728"

Now, tax at source rate is 30 %.

So, "I=70\%\ of \ 10.728=7.5096\ €"